### v1

@@ Abstract
Results: We developed ropebwt3 for efficient BWT construction and query. Ropebwt3 could index 100 assembled human genomes in 21 hours and index 7.3 terabases of commonly studied bacterial assemblies in 26 days. This was achieved using 82 gigabytes of memory at the peak without working disk space. Ropebwt3 can find maximal exact matches and inexact alignments under affine-gap penalties, and can retrieve all distinct local haplotypes matching a query sequence. It demonstrates the feasibility of full-text indexing at the terabase scale.
@@ 1 Introduction
Compressed full-text indices, such as FM-index (Ferragina and Manzini, , 2000) and r-index (Gagie et al., , 2018; Bannai et al., , 2020; Gagie et al., , 2020), provide an alternative way for fast sequence search (Navarro, , 2022). The core component of these data structures is often Burrows-Wheeler Transform (BWT; Burrows and Wheeler, 1994) which is a lossless transformation of strings. The BWT of a highly redundant string tends to group symbols in the original string into long runs and can thus be well compressed. When we supplement BWT with a data structure to efficiently compute the rank of a symbol in BWT, we can in theory count the occurrences of a string in time linear to its length. FM-index further adds a sampled suffix array to locate substrings, while r-index uses an alternative method that is more efficient for redundant strings. Both of them support compression and sequence search at the same time.
BWT-based indices have been used for read alignment (Langmead et al., , 2009; Li and Durbin, , 2009; Li et al., , 2009), de novo sequence assembly (Simpson and Durbin, , 2012) and data compression (Cox et al., , 2012). They have also emerged as competent data structures for pangenome data. Existing pangenome-focused tools (Rossi et al., , 2022; Ahmed et al., , 2021; Zakeri et al., , 2024) use prefix-free parsing for BWT construction (Boucher et al., , 2019). They require more memory than input sequences and are impractical for huge datasets. Although ropebwt2 developed by us can construct BWT in memory proportional to its compressed size and is fast for short strings (Li, , 2014), it is inefficient for chromosome-long sequences. OnlineRlbwt (Bannai et al., , 2020) has a similar limitation. grlBWT (Díaz-Domínguez and Navarro, , 2023) is likely the best algorithm for constructing the BWT of similar genomes. It reduces the peak memory at the cost of large working disk space and frequent disk input/output. In its current form, the algorithm does not support update to BWT. We would need to reconstruct the BWT from scratch when new genomes arrive. BWT construction for highly redundant sequences remains an active research area.
@@
It is also possible to identify inexact local matches. With the BWT-SW algorithm, Lam et al., (2008) simulated a suffix trie, a tree data structure, with BWT and performed sequence-to-trie alignment to find all local matches between a query string and the BWT of a large genome. We went a step further by representing the query string with its direct acyclic word graph (DAWG; Blumer et al., 1983) and performed DAWG-to-trie alignment. This is the BWA-SW algorithm (Li and Durbin, , 2010). Nonetheless, we later realized the formulation of BWA-SW had theoretical flaws and its implementation was closer to DAWG-to-DAWG alignment than to DAWG-to-tree alignment.
In this article, we will explain our solution to BWT construction and to sequence search at the terabase scale. Our contribution includes: a) a reinterpreted BWA-SW algorithm that fixes issues in our earlier work (Li and Durbin, , 2010); b) its application in finding similar haplotypes; c) an incremental in-memory BWT construction implementation that scales to large pangenome datasets.
@@ 2 Methods
In a nutshell, ropebwt3 computes the partial multi-string BWT of a subset of sequences with libsais and merges the partial BWT into the existing BWT run-length encoded as a B+-tree (Li, , 2014). It repeats this procedure until all input sequences are processed. The BWT by default includes input sequences on both strands. This enables forward-backward search (Li, , 2012) required by accelerated long MEM finding (Gagie, , 2024). Ropebwt3 also reports local alignment with affine-gap penalty using a revised BWA-SW algorithm (Li and Durbin, , 2010).
@@
$\pi(i)$ LF-mapping: $\pi(i)\triangleq S^{-1}(S(i)-1)$
@@ 2.1 Basic concepts
For $a\in\Sigma^{\prime}$, let $C_{B}(a)\triangleq|\{i:B[i]<a\}|$ be the number of symbols smaller than $a$ and ${\rm rank}_{B}(a,k)\triangleq|\{i<k:B[i]=a\}|$ be the number of $a$ before offset $k$ in $B$. We may omit subscription $B$ when we are describing one string only. The last-to-first mapping (LF mapping) $\pi$ is defined by $\pi(i)\triangleq S^{-1}(S(i)-1)$, where $S^{-1}$ is the inverse function of suffix array $S$. It can be calculated as $\pi(i)=C(B[i])+{\rm rank}(B[i],i)$. As $B[\pi(i)]$ immediately proceeds $B[i]$ on $T$, we can use $\pi$ to decode the $i$-th sequence in $B$.
@@ 2.2 BWT construction
2:     $m_{1}\leftarrow|\{k:B_{1}[k]=\$\}|$; $m_{2}\leftarrow|\{k:B_{2}[k]=\$\}|$
3:     for $k\leftarrow 0$ to $|B_{2}|-1$ do
4:         $a\leftarrow B_{2}[k]$
5:         $R(k)\leftarrow(a,C_{B_{2}}(a)+{\rm rank}_{B_{2}}(a,k))$
6:     end for
7:     for $i\leftarrow 0$ to $m_{2}-1$ do
8:         $k\leftarrow i$; $l\leftarrow m_{1}$
9:         repeat
10:              $(a,k^{\prime})\leftarrow R(k)$
11:              $R(k)\leftarrow(a,k+l)$$\triangleright$ position in the merged BWT
12:              $k\leftarrow k^{\prime}$; $l\leftarrow C_{B_{1}}(a)+{\rm rank}_{B_{1}}(a,l)$
13:         until $a=\$$
14:     end for
15:     for $(a,k)\in R$ $\triangleright$ N.B. $k$ is sorted in array $R$
16:         ${\rm insert}_{B_{1}}(a,k)$$\triangleright$ insert $a$ after $k$ symbols in $B_{1}$
17:     end for
@@
For $m$ that can be represented by a 32-bit integer and $n$ represented by a 64-bit integer, libsais will need at least $12n$ bytes to construct the suffix array. It is impractical for $n$ more than tens of billions. To construct BWT for large $n$, ropebwt3 uses libsais to build the BWT for a batch (up to 7 Gb by default) and merges it to the BWT of already processed batches (Algorithm 1). The basic idea behind the algorithm is well known (Ferragina et al., , 2010) but implementations vary (Sirén, , 2016; Oliva et al., , 2021). In ropebwt3, we encode BWT with a B+-tree (Fig. 2a). This yields $O(\log r)$ rank query (line 12) and insertion (line 16), where $r$ is the number of runs in the merged BWT. The bottleneck of the algorithm lies in rank calculation (line 7), which can be parallelized if $m_{2}>1$.
Ropebwt2 (Li, , 2014) uses the same B+-tree to encode BWT and has the same time complexity. However, it inserts sequences, not partial BWTs, into existing BWT. It cannot be efficiently parallelized for long strings. Note that independent of our earlier work, Ohno et al., (2018) also used B+-tree to encode BWT. Its implementation (Bannai et al., , 2020) is several times slower than ropebwt2 on 152 bacterial genomes from Li et al., (2024), possibly because ropebwt2 is optimized for the small DNA alphabet.
Figure 2: Examples of binary BWT encoding. (a) Encoded as a B+-tree. An internal node keeps the marginal counts of symbols in its descendants. An external node keeps the run-length encoded substring in BWT. A run length may be encoded in one, two, four or eight bytes in a scheme inspired by UTF-8. (b) Encoded as a bit-packed array. The first two rows store run-length encoded BWT interleaved with marginal counts in each block. The Elias delta encoding is used to represent run lengths. The last row is an index into BWT for fast access.
2.3 Suffix array interval and backward search
For a string $P\in\Sigma^{*}$ (i.e. not including sentinels), let ${\rm occ}(P)$ be the number of occurrences of $P$ in $T$. Define ${\rm lo}(P)$ to be the number of suffixes that lexicographically smaller than $P$ and ${\rm hi}(P)\triangleq{\rm lo}(P)+{\rm occ}(P)$. $[{\rm lo}(P),{\rm hi}(P))$ is called the suffix array interval of $P$, or SA interval in brief. An SA interval is important if there exists $P$ such that it is the SA interval of $P$. The SA interval of the empty string is $[0,n)$ where $n=|T|$. If we know the SA interval of $P$, we can calculate the SA interval of $aP$ with:
${\rm lo}(aP)$ $=$ $C(a)+{\rm rank}(a,{\rm lo}(P))$
${\rm hi}(aP)$ $=$ $C(a)+{\rm rank}(a,{\rm hi}(P))$
To calculate ${\rm occ}(P)$, we can start with $[0,n)$ and repeatedly apply the equation above from the last symbol in $P$ to the first. This procedure is called backward search.
@@ 2.4 Double-strand BWT
3:         $k^{\prime}\leftarrow k^{\prime}+\big{[}{\rm rank}(\overline{b},k+s)-{\rm rank}(\overline{b},k)\big{]}$
@@
5:     $s\leftarrow{\rm rank}(a,k+s)-{\rm rank}(a,k)$
6:     $k\leftarrow C(a)+{\rm rank}(a,k)$
@@ 2.5 Finding supermaximal exact matches
An exact match between string $T$ and $P$ is a 3-tuple $(l,t,p)$ such that $T[t,t+l)=P[p,p+l)$. A maximal exact match (MEM) is an exact match that cannot be extended in either direction. A MEM may be contained in another MEM on the pattern string $P$. For example, suppose $T={\tt GACCTCCG}$ and $P={\tt ACCT}$. MEM $(2,6,1)$ is contained in MEM $(4,1,0)$ on the pattern string. A supermaximal exact match (SMEM) is a MEM that is not contained in other MEMs on the pattern string. In the example above, only $(4,1,0)$ is a SMEM. There are usually much fewer SMEMs than MEMs. Gagie, (2024) recently proposed a new algorithm to find long SMEMs (Algorithm 3) that is faster than our earlier algorithm (Li, , 2012) especially when there are many short SMEMs to skip. Both algorithms can also find SMEMs occurring at least $s_{\min}$ times (Tatarnikov et al., , 2023).
@@
2.6 Finding inexact matches

### v2

@@ Abstract
Results: We developed ropebwt3 for efficient BWT construction and query. Ropebwt3 indexed 320 assembled human genomes in 65 hours and indexed 7.3 terabases of commonly studied bacterial assemblies in 26 days. This was achieved using up to 170 gigabytes of memory at the peak without working disk space. Ropebwt3 can find maximal exact matches and inexact alignments under affine-gap penalties, and can retrieve similar local haplotypes matching a query sequence. It demonstrates the feasibility of full-text indexing at the terabase scale.
@@ 1 Introduction
Compressed full-text indices, such as FM-index (Ferragina and Manzini, , 2000) r-index (Gagie et al., , 2018; Bannai et al., , 2020; Gagie et al., , 2020) and move index (Nishimoto and Tabei, , 2021), provide an alternative way for fast sequence search (Navarro, , 2022). The core component of these data structures is often Burrows-Wheeler Transform (BWT; Burrows and Wheeler, 1994) which is a lossless transformation of strings. The BWT of a highly redundant string tends to group symbols in the original string into long runs and can thus be well compressed. When we supplement BWT with a data structure to efficiently compute the rank of a symbol in BWT, we can in theory count the occurrences of a query string in time linear to its length. FM-index further adds a sampled suffix array to locate substrings, while r-index uses an alternative method that is more efficient for redundant strings. Both of them support compression and sequence search at the same time.
BWT-based indices have been used for read alignment (Langmead et al., , 2009; Li and Durbin, , 2009; Li et al., , 2009), de novo sequence assembly (Simpson and Durbin, , 2012), metagenome profiling (Kim et al., , 2016) and data compression (Cox et al., , 2012). They have also emerged as competent data structures for pangenome data. Existing pangenome-focused tools (Rossi et al., , 2022; Ahmed et al., , 2021; Zakeri et al., , 2024) use prefix-free parsing for BWT construction (Boucher et al., , 2019). They require more memory than input sequences and are impractical for huge datasets. Although ropebwt2 developed by us can construct BWT in memory proportional to its compressed size and is fast for short strings (Li, , 2014), it is inefficient for chromosome-long sequences. OnlineRlbwt (Bannai et al., , 2020) has a similar limitation. grlBWT (Díaz-Domínguez and Navarro, , 2023) is likely the best algorithm for constructing the BWT of similar genomes. It reduces the peak memory at the cost of large working disk space. In its current form, the algorithm does not support update to BWT. We would need to reconstruct the BWT from scratch when new genomes arrive. BWT construction for highly redundant sequences remains an active research area.
@@
It is also possible to identify inexact matches. Kucherov et al., (2016) invented search schemes to find semi-global alignment. With the BWT-SW algorithm, Lam et al., (2008) simulated a suffix trie, a tree data structure, with BWT and performed sequence-to-trie alignment to find all local matches between a query string and the BWT of a large genome. We went a step further by representing the query string with its direct acyclic word graph (DAWG; Blumer et al., 1983) and performed DAWG-to-trie alignment. This is the BWA-SW algorithm (Li and Durbin, , 2010). Nonetheless, we later realized the formulation of BWA-SW had theoretical flaws and its implementation was closer to DAWG-to-DAWG alignment than to DAWG-to-tree alignment.
In this article, we will explain our solution to BWT construction and to sequence search at the terabase scale. Our contribution includes: a) a reinterpreted BWA-SW algorithm that fixes issues in our earlier work (Li and Durbin, , 2010); b) its application in finding similar haplotypes; c) an incremental in-memory BWT construction implementation that scales to large pangenome datasets; d) a faster algorithm to find alignment positions with a standard FM-index of highly redundant strings.
@@ 2 Methods
In a nutshell, ropebwt3 computes the partial multi-string BWT of a subset of sequences with libsais (https://github.com/IlyaGrebnov/libsais) and merges the partial BWT into the existing BWT run-length encoded as a B+-tree (Li, , 2014). It repeats this procedure until all input sequences are processed. The BWT by default includes input sequences on both strands. This enables forward-backward search (Li, , 2012) required by accelerated long MEM finding (Gagie, , 2024). Ropebwt3 also reports local alignment with affine-gap penalty using a revised BWA-SW algorithm (Li and Durbin, , 2010).
@@
$\pi_{B}(a,k)$ $\pi_{B}(a,k)\triangleq C_{B}(a)+{\rm rank}_{B}(a,k)$
$\pi(k)$ LF-mapping: $\pi(k)\triangleq S^{-1}(S(k)-1)=\pi_{B}(B[k],k)$
@@ 2.1 Basic concepts
For $a\in\Sigma^{\prime}$, let $C_{B}(a)\triangleq|\{i:B[i]<a\}|$ be the number of symbols smaller than $a$ and ${\rm rank}_{B}(a,k)\triangleq|\{i<k:B[i]=a\}|$ be the number of $a$ before offset $k$ in $B$. We may omit subscription $B$ when we are describing one string only. The last-to-first mapping (LF mapping) $\pi$ is defined by $\pi(i)\triangleq S^{-1}(S(i)-1)$, where $S^{-1}$ is the inverse function of suffix array $S$. It can be calculated as $\pi(i)=\pi_{B}(B[i],i)$, where $\pi_{B}(a,i)\triangleq C(B[i])+{\rm rank}(B[i],i)$. As $B[\pi(i)]$ immediately proceeds $B[i]$ on $T$, we can use $\pi$ to decode the $i$-th sequence in $B$.
@@ 2.2 BWT construction
2:     $m_{1}\leftarrow|\{k:B_{1}[k]=\$\}|$
3:     $m_{2}\leftarrow|\{k:B_{2}[k]=\$\}|$
4:     for $k\leftarrow 0$ to $|B_{2}|-1$ do
5:         $a\leftarrow B_{2}[k]$
6:         $R(k)\leftarrow(a,\pi_{B_{2}}(a,k))$
7:     end for
8:     for $i\leftarrow 0$ to $m_{2}-1$ do
9:         $k\leftarrow i$; $l\leftarrow m_{1}$
10:         repeat
11:              $(a,k^{\prime})\leftarrow R(k)$
12:              $R(k)\leftarrow(a,k+l)$$\triangleright$ position in the merged BWT
13:              $k\leftarrow k^{\prime}$; $l\leftarrow\pi_{B_{1}}(a,l)$
14:         until $a=\$$
15:     end for
16:     for $(a,k)\in R$ $\triangleright$ N.B. $k$ is sorted in array $R$
17:         ${\rm insert}_{B_{1}}(a,k)$$\triangleright$ insert $a$ after $k$ symbols in $B_{1}$
18:     end for
19:end procedure
For $m$ that can be represented by a 32-bit integer and $n$ represented by a 64-bit integer, libsais will need at least $12n$ bytes to construct the suffix array. It is impractical for $n$ more than tens of billions. To construct BWT for large $n$, ropebwt3 uses libsais to build the BWT for a batch (up to 7 Gb by default) and merges it to the BWT of already processed batches (Algorithm 1). The basic idea behind the algorithm is well known (Ferragina et al., , 2010) but implementations vary (Sirén, , 2016; Oliva et al., , 2021). In ropebwt3, we encode BWT with a B+-tree (Fig. 2a). This yields $O(\log r)$ rank query (line 13) and insertion (line 17), where $r$ is the number of runs in the merged BWT. The bottleneck of the algorithm lies in rank calculation (line 8), which can be parallelized if $m_{2}>1$.
Ropebwt2 (Li, , 2014) uses the same B+-tree to encode BWT and has the same time complexity. However, it inserts sequences, not partial BWTs, into existing BWT. It cannot be efficiently parallelized for long strings. Note that independent of our earlier work, Ohno et al., (2018) also used a B+-tree to encode BWT. Its implementation (Bannai et al., , 2020) is several times slower than ropebwt2 on 152 bacterial genomes from Li et al., (2024), possibly because ropebwt2 is optimized for the small DNA alphabet.
Figure 2: Examples of binary BWT encoding. The alphabet is $\{{\tt A},{\tt C},{\tt G}\}$. (a) Encoded as a B+-tree. An internal node keeps the marginal counts of symbols in its descendants. An external node keeps the run-length encoded substring in BWT. A run length may be encoded in one, two, four or eight bytes in a scheme inspired by UTF-8. A B+-tree organized this way resembles the rope data structure. (b) Encoded as a bit-packed array. The first two rows store run-length encoded BWT interleaved with marginal counts in each block. The Elias delta encoding is used to represent run lengths. The last row is an index into BWT for fast access.
2.3 Suffix array interval and backward search
For a string $P\in\Sigma^{*}$ (i.e. not including sentinels), let ${\rm occ}(P)$ be the number of occurrences of $P$ in $T$. Define ${\rm lo}(P)$ to be the number of suffixes that are lexicographically smaller than $P$ and ${\rm hi}(P)\triangleq{\rm lo}(P)+{\rm occ}(P)$. $[{\rm lo}(P),{\rm hi}(P))$ is called the suffix array interval of $P$, or SA interval in brief. An SA interval is important if there exists $P$ such that it is the SA interval of $P$. The SA interval of the empty string is $[0,n)$ where $n=|T|$.
If we know the SA interval of $P$, we can calculate the SA interval of $aP$ with: ${\rm lo}(aP)=\pi_{B}(a,{\rm lo}(P))$ and ${\rm hi}(aP)=\pi_{B}(a,{\rm hi}(P))$. To calculate ${\rm occ}(P)$, we start with $[0,n)$ and repeatedly apply the equation above from the last symbol in $P$ to the first. This procedure is called backward search.
2.4 Locating with FM-index
By definition of BWT, if $i\in[{\rm lo}(P),{\rm hi}(P))$, $P=T[S(i),S(i)+|P|)$. We can thus locate an occurrence of $P$ in the original string $T$. However, suffix array $S$ takes $O(n\log n)$ in space and may be too large to store explicitly. With an FM-index, we only store $S(i)$ if and only if $i$ is a multiple of $s$ where $s$ is a positive integer controlling the sample rate. We calculate the rest of $S(i)$ using LF-mapping $\pi(\cdot)$.
A complication with multi-string BWT is that $\pi(i)$ does not point to the preceeding symbol when $B[i]=\$$. This is because sentinels are ordered during suffix sorting but are not distinguished from each other in BWT. We will have to store the rank of each string in $T$ (array $R$ in Algorithm 2). For convenience, we also keep the index of the string and the offset on the string instead of the offset on the concatenated string $T$.
To find the position of $i\in[{\rm lo}(P),{\rm hi}(P))$ in the original string, we repeatedly apply $\pi(\cdot)$ $k$ times until the position of $\pi^{k}(i)$ is stored. With $U$, $V$ and $R$ precalculated by IndexSSA in Algorithm 2, we can locate $i$ with
$\left\{\begin{array}[]{ll}(R(\pi^{k}(i)),k)&(B[\pi^{k}(i)]=\$)\\ (U(\pi^{k}(i)),V(\pi^{k}(i)+k))&(\mbox{otherwise and $\pi^{k}(i)\bmod s=0$})\\ \end{array}\right.$
where $k$ is the smallest non-negative integer such that $B[\pi^{k}(i)]=\$$ or $\pi^{k}(i)\bmod s=0$. On average, $k\approx s$, which means each locate operation triggers $s$ rank queries.
Function Locate1 in Algorithm 2 provides a faster way to find one position in SA interval $[l,h)$. A key observation is that if the interval contains a sentinel or there exists $k$ such that $l\leq ks<h$, we can immediately locate one occurrence; if not, we can apply backward search repeatedly until an SA interval $[l^{\prime},h^{\prime})$ brackets a stored suffix array value. If $T$ consists of $m^{\prime}$ identical strings, we apply $s/\min(h-l,m^{\prime})$ rounds of backward searches on average, usually much faster than the naive algorithm. Ropebwt3 implements a generalized Locate1 function that finds multiple occurrences.
Algorithm 2 Locate one hit given suffix array samples
1:procedure IndexSSA($B,s$)
2:     $A\leftarrow\emptyset$; $m\leftarrow|\{i:B[i]=\$\}|$$\triangleright$ Number of sequences
3:     for $t\leftarrow 0$ to $m-1$ $\triangleright$ Traverse all sequences
4:         $k\leftarrow t$; $l\leftarrow 0$$\triangleright$ $l$ will be the length of $t$-th sequence
5:         $\triangleright$ Iterate from the end of the sequence
6:              $k\leftarrow\pi(k)$; $l\leftarrow l+1$
7:              if $B[k]=\$$ then
8:                  $R(k)\leftarrow t$$\triangleright$ The rank of $t$-th sequence is $k$
9:              else if $k\bmod s=0$ then
10:                  $A\leftarrow A\cup\{k/s\}$
11:              end if
12:         until $B[k]=\$$
13:         for $k\in A$ do
14:              $U(k)=t$; $V(k)\leftarrow l-1-k$
15:         end for
16:     end for
17:     return $(U,V,R)$
@@ 2.4 Locating with FM-index
19:procedure Locate1($B,U,V,R,{\rm lo},{\rm hi},s$)
20:     $I\leftarrow\{({\rm lo},{\rm hi},0)\}$
21:     while $I\not=\emptyset$ do
22:         $(l,h,o)\leftarrow\mbox{largest interval in $I$}$
23:         $I\leftarrow I\setminus\{(l,h,o)\}$
24:         if $\exists k$ such that $k\cdot s\in[l,h)$ then
25:              return $(U(k),V(k)+o)$
26:         else if $\pi_{B}(\$,l)<\pi_{B}(\$,h)$ then
27:              return $(R(\pi_{B}(\$,l)),o)$
28:         end if
29:         for $a\in\Sigma$ do
30:              $I\leftarrow I\cup\{(\pi_{B}(a,l),\pi_{B}(a,h),o+1)\}$
31:         end for
32:     end while
33:end procedure
@@ 2.5 Double-strand BWT
3:         $k^{\prime}\leftarrow k^{\prime}+\big{[}\pi_{B}(\overline{b},k+s)-\pi_{B}(\overline{b},k)\big{]}$
@@
5:     $s\leftarrow\pi_{B}(a,k+s)-\pi_{B}(a,k)$
6:     $k\leftarrow\pi_{B}(a,k)$
@@ 2.6 Finding supermaximal exact matches
An exact match between strings $T$ and $P$ is a 3-tuple $(t,p,l)$ such that $T[t,t+l)=P[p,p+l)$. A maximal exact match (MEM) is an exact match that cannot be extended in either direction. A MEM may be contained in another MEM on the pattern string $P$. For example, suppose $T={\tt GACCTCCG}$ and $P={\tt ACCT}$. MEM $(5,1,2)$ is contained in MEM $(1,0,4)$ on the pattern string. A supermaximal exact match (SMEM) is a MEM that is not contained in other MEMs on the pattern string. In the example above, only $(1,0,4)$ is a SMEM. There are usually much fewer SMEMs than MEMs. Gagie, (2024) recently proposed a new algorithm to find long SMEMs (Algorithm 4) that is faster than our earlier algorithm (Li, , 2012) especially when there are many short SMEMs to skip. Both algorithms can also find SMEMs occurring at least $s_{\min}$ times (Tatarnikov et al., , 2023).
@@
2.7 Finding inexact matches with BWA-SW